Find using distributive property: 
824 x 25 ???????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given 824 × 25
= (800 + 24) × 25
= (800 + 25–1) × 25
= 800 × 25 + 25 × 25–1 × 25
= 20000 + 625–25
= 20000 + 600
= 20600

Thanks